John bought a certain number of apples at 30 cents each and he had 3 dollars left. If he bought the same
numbr of pear at 40cent eachinstead of the apples,he would be short of 1 dollar, how many apples did he buy?

Answer by ankor@dixie-net.com(22740)

John bought a certain number of apples at 30 cents each and he had 3 dollars left.
If he bought the same number of pears at 40cent each instead of the apples, he would be short of 1 dollar, how many apples did he buy?
:
Let A = amt he had to spend
let n = number of apples or pears
:
Write an equation for each transaction
A - .30n = 3
A - .40n = -1
--------------subtraction eliminates A
0 +.10n = 4
n = 4/.1
n = 40 apples (or pears)
:
:
Check this, find the Amt he had
A = .30(40) + 3
A = 12 + 3
A = $15 to speed
Check in the pear statement
15 - .40(40) =
15 - 16 = -1